OpenStudy (anonymous):

NEED HELP PLEASE!! I dont know how to find the geometric mean! Here's the problem: Find the geometric mean of 3 and 7.

OpenStudy (anonymous):

geometric mean of 2 numbers is the square root of their product. \[\sqrt{3*7}=\sqrt{21}\]

OpenStudy (anonymous):

ahh got it! thank you, this really helps!!!